A tap fills a tank 12 hours other emptied it in 14 hour if both are open how much will take time to fill it..
Answer by nerdybill(7384)

A tap fills a tank 12 hours other emptied it in 14 hour if both are open how much will take time to fill it..
.
Let x = time (hours) it takes to fill tank
then
x(1/12 - 1/14) = 1
multiplying both sides by (12)(14):
x(14 - 12) = (12)(14)
x(2) = 168
x = 84 hours